Felt Tense or Stressed Out at Work Do males or females feel more tense or stressed out at work? A survey of employed adults conducted online by a company on behalf of a research organization revealed the data in the contingency tableGender shown to the right. Complete parts (a) through (d) below No Total 355 610 Female 200225425 580 1035 Yes 255 Male Total 455

a. What is the probability that a randomly selected? person's gender is? female?

?(Type an integer or a decimal. Round to three decimal places as? needed.)

b. What is the probability that a randomly selected person feels tense or stressed out at work and is? female?

?(Type an integer or a decimal. Round to three decimal places as? needed.)

c. What is the probability that a randomly selected person feels tense or stressed out at work or is? female?

?(Type an integer or a decimal. Round to three decimal places as? needed.)

d. Explain the difference in the results in? (b) and? (c).

A.

The two characteristics in part? (b) are mutually? exclusive, while those in part? (c) are not.

B.

The joint event in part? (c) includes the possibility that just one of the two characteristics occurs.

C.

The two characteristics in part? (b) are collectively? exhaustive, while those in part? (c) are not.
Answer #1

Ans:

a)

P(female)=425/1035=0.411

b)

P(tense and female)=200/1035=0.193

c)

P(tense or female)=P(tense)+P(female)-P(tense and female)

=(455/1035)+(425/1035)-(200/1035)

=(455+425-200)/1035=0.657

d)

Option B is correct.